# Reconfigurable Intelligent Surface-Assisted Localization in OFDM Systems with Carrier Frequency Offset and Phase Noise

### v3

@@ V RIS Configuration
The partial derivatives of $\bm{\mu}[n]$ and $\bm{\mu}[n]^{*}$ with regard to all parameters to be estimated are derived in Appendix A. It can be observed that $\bm{\bar{w}}$ is only in the last term of $\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}},\forall i$, while $\bm{\bar{w}}^{\rm H}$ appears as the first term of $\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}},\forall j$, for the vector parts. Hence, the product of $\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}},\forall i$ and $\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}},\forall j$ would lead to $\mathbf{W}$. As a result, as can be seen from (51), every term of $\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})$ is linear to $\mathbf{W}$ because the operations of taking the real parts, summation, and scalar multiplication preserve linearity. Hence, $\mathbf{B}(\bm{p}_{u},\mathbf{W})$ is linear to $\mathbf{W}$. ∎
@@
Being an SDP, the problem in (V) can be solved using the interior point method in the CVX toolboxes. Thus, the convergence of the proposed RIS configuration is guaranteed. The computational complexity of the proposed RIS configuration is primarily influenced by the interior point method employed to solve (V). The worst-case complexity of the interior point method is $\mathcal{O}(\max\left\{N_{con},N_{var}\right\}^{4}\sqrt{N_{v}})$, where $N_{con}$ and $N_{var}$ are the constraint and variable numbers, respectively [39]. For the problem in (V), ${N_{con}}=N_{R}+2$ and ${N_{var}}=N_{R}^{2}+U(N+3)^{2}$. Suppose that the convergence accuracy is $\epsilon$. The complexity of solving (V) is $\mathcal{O}((N_{R}^{2}+UN^{2})^{4.5}\log\frac{1}{\epsilon})$.
@@ VI Numerical Evaluation
This section elaborates on the comprehensive simulation results conducted to assess the proposed CFO, PN and position estimation and the RIS configuration algorithms. We adopt a 3D Cartesian coordinate system, where the AN is located at $\bm{p}_{a}=[\rm 2\>m,-2\>m,1\>m]^{\top}$, and the reference point of the RIS is located at $\bm{0}_{3\times 1}{\rm\>m}$. The RIS is placed on the $yz$ plane, without loss of generality. The AOI is set to be a cube centered at $[\rm 2\>m,2\>m,0\>m]^{\top}$ with the side length of $\rm 1\>m$. The light speed is $c=3\times 10^{8}\>{\rm m/s}$, and the noise power is $\sigma^{2}=-109\;{\rm dBm}$. $N=32$ subcarriers are used in an OFDM symbol, the central frequency is $f_{c}=2.8{\rm\>GHz}$, and the bandwidth is $B=100{\rm\>MHz}$. The antenna gains $G_{a}$ and $G_{t}$ are both set to $1$. Unless otherwise specified, $N_{R}=81$ and $\sigma_{\Delta}^{2}=10^{-3}\;{\rm rad^{2}}$ are used in the simulations.
@@ VI-A Convergence of Algorithm 1
Figure 3: The convergence of the proposed Algorithm 1, where the transmit power is $P=-10{\rm\;dBm}$.
First, we illustrate the convergence of Algorithm 1. Fig. 3 illustrates the case when the transmit power is $-10{\rm\>dBm}$, which corresponds to a receive signal-to-noise ratio (SNR) of approximately $17.5{\rm\>dB}$. Specifically, Fig. 3(a) depicts the change of the objective function when operating the GD algorithm to estimate the UE position, and different curves correspond to the GD algorithm in different rounds of iteration. Fig. 3(b) focuses on the convergence of the overall algorithm, and different points on the curves represent the objective function’s values after different rounds of iteration. In general, the GD algorithm typically converges within 2,000 iterations in this setup, which converges increasingly earlier within an iteration as the overall algorithm runs. The overall algorithm generally converges within 55 iterations.
TABLE II: Average number of iterations for Algorithm 1
@@
Avg. No. of Iterations 19.45 50.59 84.31 134.95
Further, Table II shows the average number of iterations under different transmit powers for Algorithm 1. In this simulation, 100 tests are conducted, and $N_{o}$ and $N_{i}$ are set to be sufficiently large. The results indicate that the required iteration number increases as the transmit power increases.
VI-B Performance of RIS Configuration
@@
Figure 4: The PEB within AOI for various RIS configuration strategies.
@@
Fig. 4 indicates that the proposed optimization algorithm improves the PEB by approximately two orders of magnitude by comparison with the random RIS phase shifts, and by approximately one order of magnitude compared to the situation where there are no hardware impairments [27]. This fully demonstrates the significance of considering the CFO and PN in the localization of the UE, as well as the effectiveness, progressiveness, and superiority of the proposed RIS configuration algorithm. Additionally, observed from Fig. 4, the optimized RIS phase shifts not only improve the localization accuracy within the AOI but also significantly improve localization performance in its surrounding areas.
@@ VI-C Estimation Accuracy
Figure 5: The CDF of RMSE and PEB with different $N_{R}$ and transmit powers.
Then, we concentrate on the CFO, PN, and UE position estimation accuracy. We first illustrate the overall localization accuracy. Fig. 5 illustrates the cumulative distribution functions (CDFs) of RMSE and PEB for 500 random testing positions, under different transmit powers and different numbers of RIS elements. It is observed that increasing the RIS size and the transmit power results in significant enhancement of the overall localization accuracy. Furthermore, when the RIS size is $N_{R}=49$, the CDF of the RMSE deviates significantly from the CDF of the PEB, and does not achieve the theoretical localization accuracy. However, when the RIS size is increased to $N_{R}=121$, the CDF of the RMSE curve aligns more closely with the CDF of the PEB. This phenomenon can also be observed in Fig. 8.
Next, we select a fixed testing position with the coordinates $\rm[1.50\>m,2.15\>m,0.45\>m]^{\top}$ to investigate the impact of the transmit power $P$, the RIS size $N_{R}$, and the PN variance $\sigma_{\Delta}^{2}$ on the CFO, PN, and UE position estimation accuracy.
Figure 6: The RMSE of $\bm{p}_{u}$ estimation under random and optimized RIS phase shifts.
@@
In Figs. 6 and 7, we assess the effect of the optimized RIS phase shifts on the performance of estimation, compared to the random setup. From Fig. 6, the localization performance with random RIS phase shifts does not experience a significant decline as the transmit power increases. At the other extreme, the optimized RIS phase shifts provide an improvement of approximately two orders of magnitude in localization accuracy. This demonstrates the crucial role of optimizing the RIS phase shifts in localization and underscores the effectiveness of the proposed RIS configuration optimization algorithm. On the other hand, Fig. 7 shows that compared to the optimized RIS phase shifts, using random RIS phase shifts yields even lower joint MSE of CFO and PN, although the curve does not experience any significant decline as the transmit power increases in this case. This occurs because we extract the PEB part of HCRLB and set it as the RIS optimization objective. Thus, the estimation performance of the CFO and PN parameters is ignored.
Figure 8: The RMSE of $\bm{p}_{u}$ estimation at $N_{R}=49$, $N_{R}=81$, and $N_{R}=121$.
Figure 9: ${\rm MSE}_{\phi,\bm{\theta}}$ at $N_{R}=49$, $N_{R}=81$, and $N_{R}=121$.
Figs. 8 and 9 depict the curves of the RMSE for UE localization and MSE for the estimation of $\underline{\bm{\gamma}}$, respectively, under different transmit powers, where $N_{R}$ is 49, 81, or 121. It is observed that increasing the transmit power and/or the RIS element number results in gains in all parameters to be estimated. However, as the transmit SNR increases, the RMSE and MSE do not decrease as sharply as HCRLB because of the approximate PN vector, as well as the Taylor series approximation considered when deriving closed-form expressions for the CFO and PN.
Figure 10: ${\rm RMSE}_{u}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$.
@@
Figs. 10 and 11 plot the RMSE for localization and MSE for estimation of $\underline{\bm{\gamma}}$, respectively, when $N_{R}$ is set to 81, where the PN variance is $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ or $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$. The results demonstrate that decreasing the PN covariance leads to significant improvements in the estimation performance of the CFO, PN, and UE position. Another important finding is that a smaller PN variance would make the RMSE and MSE curves more closely aligned with the HCRLB curves because of the more accurately approximated parameters when updating the CFO and PN parameters.
@@
Figure 12: Benchmarking the performance of the proposed algorithms against the state-of-the-art approach in [27].
@@
Last but not least, we compare the proposed overall algorithm of the CFO, PN, and UE position estimation and the RIS configuration, with the benchmark algorithm developed in [27]. It is shown in Fig. 12 that, after optimization, the theoretical localization accuracy of our proposed algorithm outperforms the benchmark, which confirms the conclusion drawn in Fig. 4(b). As for the localization RMSE, the benchmark shows a large gap with its corresponding PEB, and its performance shows no demonstrable correlation with the transmit power because the estimation algorithm totally overlooked the CFO and PN. In comparison, the proposed algorithm significantly improves the localization accuracy by about one order of magnitude under relatively low transmit powers of below $-10{\rm\>dBm}$, and by two orders of magnitude under high transmit powers of over $10{\rm\>dBm}$.

### v4

@@ V RIS Configuration
The partial derivatives of $\bm{\mu}[n]$ and $\bm{\mu}[n]^{*}$ with regard to all parameters to be estimated are derived in Appendix A. It can be observed that $\bm{\bar{w}}$ is only in the last term of $\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}},\forall i$, while $\bm{\bar{w}}^{\rm H}$ appears as the first term of $\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}},\forall j$, for the vector parts. Hence, the product of $\frac{\partial\bm{\mu}[n]^{*}}{\partial\lambda_{i}},\forall i$ and $\frac{\partial\bm{\mu}[n]}{\partial\lambda_{j}},\forall j$ would lead to $\mathbf{W}$. As a result, as can be seen from (55), every term of $\bm{\mathrm{FIM}}(\bm{y};\bm{\lambda})$ is linear to $\mathbf{W}$ because the operations of taking the real parts, summation, and scalar multiplication preserve linearity. Hence, $\mathbf{B}(\bm{p}_{\color[rgb]{0,0,0}s},\mathbf{W})$ is linear to $\mathbf{W}$. ∎
@@
Being an SDP, the problem in (V) can be solved using the interior point method in the CVX toolboxes. Thus, the convergence of the proposed RIS configuration is guaranteed. The computational complexity of the proposed RIS configuration is primarily influenced by the interior point method employed to solve (V). The worst-case complexity of the interior point method is $\mathcal{O}(\max\left\{N_{\color[rgb]{0,0,0}\text{con}},N_{\color[rgb]{0,0,0}\text{var}}\right\}^{4}\sqrt{N_{\color[rgb]{0,0,0}\text{var}}}{\color[rgb]{0,0,0}\log\frac{1}{\epsilon}})$, where $N_{\color[rgb]{0,0,0}\text{con}}$ and $N_{\color[rgb]{0,0,0}\text{var}}$ are the constraint and variable numbers, respectively, and $\epsilon$ is the convergence accuracy [48]. For the problem in (V), ${N_{\color[rgb]{0,0,0}\text{con}}}={\color[rgb]{0,0,0}N_{\text{R}}}+2$ and ${N_{\color[rgb]{0,0,0}\text{var}}}={\color[rgb]{0,0,0}N_{\text{R}}}^{2}+{\color[rgb]{0,0,0}S}(N+3)^{2}$. Suppose that the convergence accuracy is $\epsilon$. The complexity of solving (V) is $\mathcal{O}(({\color[rgb]{0,0,0}N_{\text{R}}}^{2}+{\color[rgb]{0,0,0}S}N^{2})^{4.5}\log\frac{1}{\epsilon})$. In comparison, the complexity of the state of the art [26] is $\mathcal{O}((N_{\text{R}}^{2}+9{\color[rgb]{0,0,0}S})^{4.5}\log\frac{1}{\epsilon})$. While incurring a slightly higher complexity, our method improves the PEB by threefold over the benchmark, due to our consideration of CFO and PN, as verified numerically in Section VI-B.
@@ VI Numerical Evaluation
This section elaborates on the comprehensive simulation results conducted to assess the proposed CFO, PN and UE position estimation and the RIS configuration algorithms. We adopt a 3D Cartesian coordinate system, where the AN is located at ${\color[rgb]{0,0,0}\bm{p}_{\text{a}}}=[\rm 2\>m,-2\>m,1\>m]^{\top}$, and the reference point of the RIS is located at $\bm{0}_{3\times 1}{\rm\>m}$. The RIS is placed on the $yz$ plane, without loss of generality. The AOI is set to be a cube centered at $[\rm 2\>m,2\>m,0\>m]^{\top}$ with the side length of $\rm 1\>m$. The light speed is $c=3\times 10^{8}\>{\rm m/s}$, and the noise power is $\sigma^{2}=-109\;{\rm dBm}$. $N=32$ subcarriers are used in an OFDM symbol, the central frequency is ${\color[rgb]{0,0,0}f_{\text{c}}}=2.8{\rm\>GHz}$, and the bandwidth is $B=100{\rm\>MHz}$. The antenna gains $\color[rgb]{0,0,0}G_{\text{a}}$ and $\color[rgb]{0,0,0}G_{\text{t}}$ are both set to $1$. Unless otherwise specified, ${\color[rgb]{0,0,0}N_{\text{R}}}=81$ and $\sigma_{\Delta}^{2}=10^{-3}\;{\rm rad^{2}}$ are used in the simulations.
@@ VI-A Convergence of Algorithm 1
Figure 3: The convergence of the proposed Algorithm 1, where the transmit power is $P=10{\rm\;dBm}$.
First, we demonstrate the convergence of Algorithm 1. Fig. 3 illustrates the case with the transmit power of $-20{\rm\>dBm}$, which corresponds to a receive SNR of approximately $37.5{\rm\>dB}$. Specifically, Fig. 3(a) depicts the change of the objective function when operating the GD algorithm to estimate the UE position. The $k$-th curve ($k=0,1,\cdots$) from top in Fig. 3(a) shows the convergence of the GD algorithm in the $k$-th iteration of Algorithm 1. The GD algorithm demonstrates good convergence and effectively reduces the objective function. Fig. 3(b) focuses on the convergence of the overall algorithm, and different points on the curves represent the objective function’s values after different rounds of iteration. The GD algorithm generally converges within 3,000 iterations in this setup, and converges increasingly faster within an iteration. The overall algorithm generally converges within 150 iterations. Moreover, Table II shows the average number of iterations of Algorithm 1 and the RMSE of UE localization under different transmit powers, $N_{\color[rgb]{0,0,0}\text{O}}$ and $N_{\color[rgb]{0,0,0}\text{I}}$ are set to be sufficiently large, and 100 tests are conducted. The results indicate that as the transmit power increases, the number of iterations required decreases while the localization accuracy improves under the proposed algorithm.
VI-B Performance of RIS Configuration
TABLE II: Average number of iterations and UE localization RMSE for Algorithm 1, where $\epsilon_{\text{o}}=10^{-7}\times 10^{\rm Round(\log\mathcal{L})}$.
@@ VI-B Performance of RIS Configuration
Avg. No. of Iterations 1238.60 327.90 195.16 137.74
RMSE (m) $0.1590$ $0.0503$ $0.0197$ $0.0090$
@@
Figure 4: The logarithm of PEB within AOI for various RIS configuration strategies.
@@
Fig. 4 indicates that the proposed optimization algorithm improves the PEB by approximately two orders of magnitude compared to the random RIS phase shifts, and by more than threefold compared to the situation overlooking hardware impairments [26]. This demonstrates the significance of considering the CFO and PN in the localization of the UE, as well as the effectiveness, progressiveness, and superiority of the proposed RIS configuration algorithm. Additionally, observed from Fig. 4, the optimized RIS phase shifts not only improve the localization accuracy within the AOI but also significantly improve localization performance in its surrounding areas.
(a) Different RSR volumes
(b) Different sampling strategies
Figure 5: The effects of different RSR volumes and random sampling strategies on the CDF of PEB within AOI.
Furthermore, we explore the effects of the size of the random sampling region (RSR), and the strategy of random sampling on localization. Note that the RSR with volume $V$, coincides with the AOI by default in the RIS configuration algorithm. Fig. 5(a) shows the impact of different RSR strategies on the localization performance. By setting the center of the RSR to the center of the AOI, it is observed that as $V$ changes, the CDF of PEB does not vary significantly. This is because the proposed RIS configuration algorithm can improve the localization accuracy within the encompassed region of the RSR, as already revealed in Fig. 4.
Fig. 5(b) shows the effect of different random sampling strategies on the CDF of PEB within AOI. Uniform and Gaussian sampling are considered. The strategy of random sampling significantly affects the localization performance, and uniform sampling outperforms Gaussian sampling. Moreover, random sampling affects the performance of the RIS configuration algorithm and the localization accuracy more significantly than the RSR strategy, as also shown in Fig. 5(a).
@@ VI-C Estimation Accuracy
Figure 6: The CDF of RMSE and PEB with different ${\color[rgb]{0,0,0}N_{\text{R}}}$ and transmit powers.
Then, we concentrate on the CFO, PN, and UE position estimation accuracy. We first illustrate the overall localization accuracy. Fig. 6 illustrates the cumulative distribution functions (CDFs) of RMSE and PEB for 500 random testing positions, under different transmit powers and different numbers of RIS elements. It is observed that increasing the RIS size and the transmit power results in significant enhancement of the overall localization accuracy. Furthermore, when the RIS size is ${\color[rgb]{0,0,0}N_{\text{R}}}=49$, the CDF of the RMSE deviates significantly from the CDF of the PEB, and does not achieve the theoretical localization accuracy. However, when the RIS size is increased to ${\color[rgb]{0,0,0}N_{\text{R}}}=121$, the CDF of the RMSE curve aligns more closely with the CDF of the PEB. This phenomenon can also be observed in Fig. 9.
Next, we select a fixed testing position with the coordinates $\rm[1.50\>m,2.15\>m,0.45\>m]^{\top}$ to investigate the impact of the transmit power $P$, the RIS size ${\color[rgb]{0,0,0}N_{\text{R}}}$, and the PN variance $\sigma_{\Delta}^{2}$ on the CFO, PN, and UE position estimation accuracy.
Figure 7: The RMSE of $\bm{p}_{\color[rgb]{0,0,0}\text{u}}$ estimation under random and optimized RIS phase shifts.
@@
In Figs. 7 and 8, we assess the effect of the optimized RIS phase shifts on the performance of estimation, compared to the random setup. From Fig. 7, the localization performance with random RIS phase shifts does not experience a significant decline as the transmit power increases. At the other extreme, the optimized RIS phase shifts provide an improvement of approximately two orders of magnitude in localization accuracy. This demonstrates the crucial role of optimizing the RIS phase shifts in localization and underscores the effectiveness of the proposed RIS configuration algorithm. On the other hand, Fig. 8 shows that compared to the optimized RIS phase shifts, using random RIS phase shifts yields even lower joint MSE of CFO and PN, although the curve does not experience any significant decline as the transmit power increases in this case. This occurs because we extract the PEB part of HCRLB and set it as the RIS optimization objective. Thus, the estimation performance of the CFO and PN parameters is ignored.
Figure 9: The RMSE of $\bm{p}_{\color[rgb]{0,0,0}\text{u}}$ estimation at ${\color[rgb]{0,0,0}N_{\text{R}}}=49$, ${\color[rgb]{0,0,0}N_{\text{R}}}=81$, and ${\color[rgb]{0,0,0}N_{\text{R}}}=121$.
Figure 10: ${\rm MSE}_{\phi,\bm{\theta}}$ at ${\color[rgb]{0,0,0}N_{\text{R}}}=49$, ${\color[rgb]{0,0,0}N_{\text{R}}}=81$, and ${\color[rgb]{0,0,0}N_{\text{R}}}=121$.
Figs. 9 and 10 depict the curves of the RMSE for UE localization and MSE for the estimation of $\underline{\bm{\gamma}}$, respectively, under different transmit powers, where ${\color[rgb]{0,0,0}N_{\text{R}}}$ is set to 49, 81, or 121. It is observed that increasing the transmit power and/or the RIS element number results in gains in all parameters to be estimated. However, as the transmit SNR increases, the RMSE and MSE do not decrease as sharply as HCRLB, because of the approximate PN vector, as well as the Taylor series approximation considered when deriving closed-form expressions for the CFO and PN.
Figure 11: ${\rm RMSE}_{\text{u}}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$.
@@
Figs. 11 and 12 plot the RMSE for localization and MSE for estimation of $\underline{\bm{\gamma}}$, respectively, when ${\color[rgb]{0,0,0}N_{\text{R}}}$ is set to 81, where the PN variance is $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ or $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$ [49, 50, 51, 52]. $\sigma_{\Delta}^{2}=10^{-3}\>{\rm rad^{2}}$ is a considerably high PN variance [49]. The results demonstrate that decreasing the PN covariance leads to significant improvements in the estimation performance of the CFO, PN, and UE position. Another important finding is that a smaller PN variance would make the RMSE and MSE curves more closely aligned with the HCRLB curves because of the more accurately approximated parameters when updating the CFO and PN parameters.
@@
Figure 13: Benchmarking the performance of the proposed algorithms against the state-of-the-art approach in [26], as well as the case with perfectly known CFO and PN.
@@
Furthermore, we compare the proposed overall algorithm of the CFO, PN, and UE position estimation and the RIS configuration, with the benchmark algorithm developed in [26]. It is shown in Fig. 13 that, after optimization, the theoretical localization accuracy of our proposed algorithm outperforms the benchmark, which confirms the conclusion drawn in Fig. 4(b). As for the localization RMSE, the benchmark shows a large gap with its corresponding PEB, and its performance shows no demonstrable correlation with the transmit power because the estimation algorithm totally overlooked the CFO and PN. In comparison, the proposed algorithm significantly improves the localization accuracy by approximately one order of magnitude under relatively low transmit powers of below $-10{\rm\>dBm}$, and by two orders of magnitude under high transmit powers of over $10{\rm\>dBm}$.
Fig. 13 also depicts the localization accuracy when the CFO and PN values are perfectly known. Compared to the proposed algorithm and the benchmark, effective mitigation of the CFO and PN effects can play a crucial role in improving localization accuracy. Furthermore, the utilization of our algorithm results in localization accuracy that is very close to the optimal value with known CFO and PN.
Figure 14: The impact of the number of antenna elements at the UE on the PEB under different settings of $P$ and $\sigma_{\Delta}^{2}$.
Figure 15: The impact of the number of antenna elements at the UE on the HCRLB of CFO and PN under different settings of $\sigma_{\Delta}^{2}$.
In Figs. 14 and 15, we investigate the impact of the number of antennas at the UE, $M$, on UE localization accuracy. The antennas are arranged in a uniform linear array along the $y$-axis, with spacing of $d$. Combining signals received by all antennas, we have $\bm{y}_{\text{com}}=\bm{\mu}_{\text{com}}+\bm{v}_{\text{com}}$, where $\bm{y}_{\text{com}}=\left[\bm{y}_{0}^{\top},\bm{y}_{1}^{\top},\cdots,\bm{y}_{M-1}^{\top}\right]^{\top}$, $\bm{\mu}_{\text{com}}=\left[\bm{\mu}_{0}^{\top},\bm{\mu}_{1}^{\top},\cdots,\bm{\mu}_{M-1}^{\top}\right]^{\top}$, and $\bm{v}_{\text{com}}=\left[\bm{v}_{0}^{\top},\bm{v}_{1}^{\top},\cdots,\bm{v}_{M-1}^{\top}\right]^{\top}$. $\bm{y}_{m}$, $\bm{\mu}_{m}$, and $\bm{v}_{m}$, $m=0,1,\cdots,M-1$, represent the received signal, noise-free received signal, and combined noise at the $m$-th antenna of the UE, respectively. $\bm{\mu}_{m}$ can be further expressed as $\bm{\mu}_{m}=\sqrt{P}\bm{\Lambda}_{\phi}\bm{\Lambda}_{\bm{\theta}_{m}}\mathbf{S}^{\top}\bm{h}_{m}$, where $\bm{h}_{m}$ can be obtained by substituting the coordinates of the UE’s $m$-th antenna into (7). $\bm{\theta}_{m},\,m=0,1,\cdots,M-1$, is the PN vector at the UE’s $m$-th antenna and is independent and identically distributed [53].
Fig. 14 plots the PEB with the increase in $M$. The impacts of the PN variance $\sigma_{\Delta}^{2}$ and the RIS size $N_{\text{R}}$ are considered. It is observed that as $M$ increases, the analytical localization accuracy error bound, i.e., PEB, decreases significantly. This indicates that increasing the number of UE antennas, like increasing the transmit power and the number of RIS elements, is effective in improving localization accuracy. Under our experimental settings, compared to a single-antenna UE, adding 4 receive antennas yields a greater localization performance gain than adding 40 RIS elements; adding 8 receive antennas yields a localization gain greater than increasing the transmit power by 10 dBm.
Fig. 15 depicts the respective HCRLBs of the combined PN vector, $\bm{\theta}_{\rm com}\triangleq\left[\bm{\theta}_{0},\bm{\theta}_{1},\cdots,\bm{\theta}_{M-1}\right]$, the PN vector at the 00-th UE, $\bm{\theta}_{0}$, and the CFO, $\phi$, under different $\sigma_{\Delta}^{2}$ settings. It is observed that the HCRLBs of $\bm{\theta}_{0}$ and $\phi$ decrease as $M$ increases, due to the additional information introduced by increasing the antennas at the UE. The HCRLB of $\phi$ decreases faster than that of $\bm{\theta}_{0}$, because the rise of $M$ increases observations of $\phi$ but does not increase observations of $\bm{\theta}_{0}$. Nevertheless, the improved estimation accuracy of $\phi$ and UE position indirectly improves the estimation accuracy of $\bm{\theta}_{0}$. By contrast, the HCRLB of the combined PN vector increases with $M$, as the increase of $M$ results in more PN vectors that require estimation. Even though the estimation of the PN on each individual antenna of the UE becomes more accurate, considering all PN vectors, the overall estimation accuracy deteriorates.
VI-D IEEE 802.11n Setup
(a) Random RIS phase shift
(b) Proposed optimization algorithm
Figure 16: The logarithm of PEB within AOI for various RIS configuration strategies under standard IEEE 802.11n settings.
Figure 17: ${\rm RMSE}_{\text{u}}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$ under standard IEEE 802.11n settings.
Figure 18: ${\rm MSE}_{\phi,\bm{\theta}}$ at $\sigma_{\Delta}^{2}=10^{-3}{\rm\>rad^{2}}$ and $\sigma_{\Delta}^{2}=10^{-4}{\rm\>rad^{2}}$ under standard IEEE 802.11n settings.
Following the IEEE 802.11n (Wi-Fi 4) [54] settings, the carrier frequency $f_{\color[rgb]{0,0,0}\text{c}}$ is set to $2.4\>{\rm GHz}$, $N=64$ subcarriers with $8$ null subcarriers are utilized, and the bandwidth $B$ is $20\>{\rm MHz}$. Fig. 16 plots the logarithm of PEB near the AOI under both random and optimized RIS configurations. The proposed algorithm improves the PEB by approximately two orders of magnitude, compared to the random RIS phase shifts. Figs. 17 and 18 plot the RMSE and joint MSE of CFO and PN under different $\sigma_{\Delta}^{2}$ settings, respectively. It is observed that decreasing the PN covariance improves the estimation performance of the CFO, PN, and UE position significantly. Overall, Figs. 16, 17, and 18 provide consistent conclusions with those drawn in Figs. 4, 11, and 12, demonstrating the ability of our algorithms to be extended to practical systems.